Hi I am in school and trying to figure out how many points my teacher is offering.
To start out(1) we have 9 quizzes which are worth 10pts each, 1 of the 10 pt quizzes are dropped. Section 1 is worth 35% of our grade(2) We have a midterm paper wort 30% of our grade. (3) We have a final worth 35% of our grade. The teacher has not told us what the point system is and did not tell us out of how many points the 2nd and 3rd sections are worth. Is there a way to figure this out?

Answer by bucky(2189)

This is a little tricky ... depending on how your teacher grades. So I'm going to make a few
assumptions.
.
I'm going to begin by assuming that your teacher wants your final grade to be based
on 100 points. Therefore, at the end of the term your teacher will come up with a numerical
score for your performance. If you do everything perfectly, you will get 100 as your term
grade.
.
Since your final exam counts 35% of your grade, your final exam could contribute up to
35 points to your term grade. For example, if you score 100 on your final exam, you will
get 35 points for your term grade. But if you score 50 points on your final you will get only
17.5 points for your term grade. Following this logic, if your teacher scores your final
exam on the basis of 100 points is perfect and 0 points means you got nothing correct, you
can multiply your exam score by 35% (or its decimal equivalent of 0.35) and find out how many
points you added to your term grade.
.
Next, your term paper counts 30% of your final grade. I'm going to assume that your
teacher also scores your term paper on a 100 point scale instead of grading it as a
letter grade ... A, B, C, D, or F. If you get a 90 on your term paper then you would multiply
90 times 30% (or its decimal equivalent of 0.30) to find how many points you get for your
term grade. (90 * 0.30 = 27.0 points.)
.
Finally, the tricky part ... Section 1. This section must contribute the remaining
35% to your 100-point term grade. Therefore, Section 1 is worth 35 points, and since there
are 8 quiz grades that count, each of the 8 quiz grades are worth 4.375 points to your
term grade. (The 4.375 points comes from dividing 35 points by 8.) And since each quiz
has 10 points, each point you score on a quiz that you count contributes 0.4375 points to your
term score.
.
How about doing an example.
.
The 8 quiz scores that you keep are 10, 5, 8, 8, 9, 10, 7, and 9. Multiply each of these
scores by 0.4375 and add them up to get how much they contribute to your final grade for
the term. Here are the results:
.
10 * 0.4375 = 4.375
5 * 0.4375 = 2.1875
8 * 0.4375 = 3.500
8 * 0.4375 = 3.500
9 * 0.4375 = 3.9375
10 * 0.4375 = 4.375
7 * 0.4375 = 3.0625
9 * 0.4375 = 3.9375
.
Now add the 8 products ... 4.375 + 2.1875 + 3.5 + 3.5 + 3.9375 + 4.375 + 3.0625 + 3.9375
to get that the part of your term score that you get from quizzes. In this example
that sum is 28.875 points.
.
Next suppose that you get a score of 89 on your mid-term paper. Multiply this by 0.30 and
you get 26.7 points toward your term score.
.
Finally, suppose that you get a score of 96 on your final exam. Multiply that score by
0.35 and you get 33.6 points for your term grade.
.
Now to get your term grade add all the points from the three sections ... in this example
the term grade will be the total of 28.875 + 26.7 + 33.6. The sum of these scores is
89.175 and that is the score that will be used as your grade for the term.
.
Notice that if you scored 10 on every one of the 8 quizzes, you would get 4.375 times 8 or 35
points toward your final grade. Then if you got 100 on your mid-term paper you would get
100 * 0.30 or 30 points of your final grade. Finally, if you wrote 100 on your final exam
you would get 100 * 0.35 or 35 points toward your final grade. The total of these three
contributing sections is 35 + 30 + 35 or 100 points ... just as you would expect it to be.
.
Hope this helps you to understand your teacher's idea of how to compute your final grade.
.